4 NH_3 + 5O_2 4NO + 6 H_2O How many grams of NH_
Answer #1

Q1.

calculate mol of O2:

mol = mass/MW = 43.4/32 = 1.35625 mol of O2

ratio is:

5 mol of O" --> 4 mol of NH3

then

1.35625* 4/5 = 1.085 mol of NH3 reqruied

so:

mass = mol*MW = 1.085*17 = 18.445 g of NH3

Q2.

m = 23.4 g is theoretical...

actual = 15.6%

yield % = actual/theoretical * 100%

yield % = 15.6/23.4 *100

% Yield = 66.66 %

Q3.

M = mol/V

mol of NaOH = mass/MW = 12.6/39.997 = 0.31502 mol of NaOH

V = 625 mL = 0.625 L

M = 0.31502/0.625

M = 0.504032 M

Q4.

15% g/mL so:

350*10^2 = 350 mL --> 0.15*350 = 52.5 g of glucose

Q5.

% mm = mass of Pt / total mass * 100%

% mm = 6.5 / (12+6.5) * 100 = 35.13 %